# Quantum computing and the stable set problem

### v1

@@ 4.2 Analysis of samples
Let us return to Example 5. For a subset of vertices $X\subseteq V$ with $|X|=30$ and $|E(G[X])|=5$, Proposition 4 assures us that we can always extract a stable set of size at least $25$ from it. However, it may occur, for instance, that one of the vertices has a degree of $5$, and by removing it, we obtain a stable set of size $29$. This property is generalized with the annihilation number of the graph, a graph parameter that is also an upper bound on the stability number of a graph, as shown in [30].
@@
The original paper details the justification for this upper bound, but the concept is rather straightforward. Let degrees of vertices be defined as above. In each iteration, we remove one vertex, and we repeat this until we have a stable set. To accelerate the formation of a stable set, we aim to remove the maximum number of edges in each iteration. Consequently, we prioritize removing the vertex with the highest degree in the graph, specifically vertex $v_{n}$. Removing the vertex of degree $d_{n}$, the sum of all degrees drops by $2d_{n}$. Here we assume the best case scenario, which is that this reduction occurs in vertices $v_{i},\ldots,v_{i+d_{n}}$, where $v_{i}$ represents the first non-zero degree vertex. By preserving the highest degree vertices for as long as possible, we accelerate the attainment of a stable set. When we remove enough vertices, the total number of edges will drop to zero, leaving the remaining vertices as the upper bound for the stability number of $G$.
@@
Altogether, combining Proposition 4, Theorem 7, and Observation 8, we develop the method for enhancing the solutions returned by a quantum annealer. In the following text, we provide a detailed explanation of this method, which we call the post-processing procedure.
@@ 4.3 Post-processing procedure
Let $S=S(G,\beta)$ denote the solver we use to obtain the initial array of solutions $[X]=[X_{1},\ldots,X_{k}]$, and let $S_{post}=S(G[X_{i}],\beta_{post})$ denote the solver for recomputing the stability number of a sample $X_{i}$. The post-processing steps are as follows: first, we take a look at $X_{1}$, which has the lowest cost, and define the current best solution as $\operatorname{\textit{best}}=|X_{1}|-|E(G[X_{1}])|$. If $X_{1}$ is a stable set, then $|E(G[X_{1}])|=0$, and $|X_{1}|=-\text{cost}(X_{1},\beta)$. Otherwise, from Proposition 4, we know that we can extract a stable set of size $|X_{1}|-|E(G[X_{1}])|$ from $X_{1}$.
@@ 5 Dealing with large instances
As mentioned in Section 1.2, some graphs are too large to be directly embedded on the QPU. In such cases, we can employ the method of CH-partitioning and divide a graph into multiple smaller subgraphs, which can then be embedded on the QPU. This process leverages a related problem, namely the maximum clique problem, but, as we will observe, the conversion between the maximum clique is trivial. In Section 5.1, we describe the CH-partitioning method, and in Section 5.2, we derive a simplified version we will use in our work. Furthermore, in Section 5.3, we discuss bounding procedures that enable us to reduce the number of partitions we have to consider and present an algorithmic framework that combines the proposed simplified partitioning with the post-processing procedure.
@@ 5.1 CH-partitioning
Assume we use the presented CH-partitioning method in the following way. Let $G=(V,E)$ be a graph, where the vertices $V$ are ordered in some ordering, and set $C_{i}=\{v_{i}\}$ for all $i\in\{1,\ldots,n\}$. Then $H_{i}=N(v_{i})\setminus\{v_{i}\}=N(v_{i})$ for all $i\in\{1,\ldots,n\}$. Altogether, this CH-partition consists of $n$ pairs $(C_{i},H_{i})$ and its cost is $\Delta(G)+1$. This is the CH-partition with the lowest cost function if we enumerate over all $s$ since each partition has to contain in each set $C_{i}$ at least one vertex and all of the neighbors of $C_{i}$.
An advantage of this approach is that the computation of the CH-partitioning itself is not computationally intensive. However, this method also has a disadvantage. Let us assume that $K$ is a maximum clique such that $|K|=k$. Then, $k$ partitions contain the same maximum clique. Thus, we perform computations on $k$ sets containing the same maximum clique. To overcome this obstacle, we propose a simplified version of the CH-partitioning.
@@ 5.2 Simple CH-partitioning
We define simple CH-partitioning as follows.
@@
Finally, we note that using simple partitioning greatly reduces the cost of partitioning. We consider some instances from the literature that cannot be directly embedded on QPU and compare the costs of regular CH-partitioning for $s=n$ with the costs for the proposed simplified version, where we sort the vertices based on their degrees in increasing order. The results are presented in Table 1.
@@ 5.3 Bounding the partitions
Furthermore, we again use the annihilation number, as defined in Definition 6. Recall that the annihilation number $a(G)$ of a graph $G$ is the largest integer $k$ such that there exist $k$ different vertices in $G$ with the degree sum at most $m$, and represents an upper bound on the stability number of $G$, see Theorem 7. As previously mentioned, bounding the stability number with the annihilation number is suitable for sparse instances. However, as the number of partitions approaches $n$, and according to the definition of simple CH-partitioning, the partitions will become sparser since we remove all previously considered vertices from the neighborhoods in the complements of the partitions. Hence, we can use the annihilation number of partitions as a reduction strategy. Again, if the annihilation number of some partition is at most the size of the largest stable set in the other partitions, we can disregard it. We also utilize post-processing on every solution we calculate to remove excess edges and to maybe improve the solution even further.
@@
Finally, we note that in our computational implementation of Algorithm 2, we sort the vertices based on their degrees in increasing order. Furthermore, we use solvers $S$, $S_{post}$ and penalty terms $\beta$, $\beta_{post}$ as explained in Section 4.3.
@@ 6 Computational results
All computations presented in this work were conducted on an Intel(R) Core(TM) i7-9750H with 16GB of RAM. Additionally, we used two cloud solvers from the D-Wave Leap platform: the QPU solver Advantage_system4.1 and the hybrid solver hybrid_binary_quadratic_model_version2. We also utilized a classical, locally run simulated annealing solver, which is part of the D-Wave’s Python library we employed for interacting with the QPU.
@@ 6.2.1 Results for smaller instances
A potential drawback of our method of post-processing samples when computations are carried out with $\beta<\frac{1}{2}$ is that they are more computationally expensive compared to computations performed with $\beta\geq\frac{1}{2}$. Increased computational cost arises from the necessity to post-process a larger quantity of larger samples. As an illustration, for C125.9 with $\beta=\frac{1}{10}$ we had to recompute stability numbers on all $1000$ samples. Recall that we recalculate the stability numbers of samples by considering their connected components. To explore the structure of the samples and evaluate the complexity of these recalculations, the cardinalities of the largest components considered are presented in Table 3.
@@
Instance $n$ $\beta=1/10$ $\beta=1/8$ $\beta=1/6$ $\beta=1/4$ $\beta=1/2$ $\beta=1$ $\beta=10$ $\beta=100$ C125.9 125 67 60 47 18 3 0 0 2 dsjc125.5 125 26 26 19 19 21 14 0 15 dsjc125.9 125 66 59 45 24 0 0 0 0 evil_chv12x10 120 71 50 31 12 6 4 0 0 evil_myc5x24 120 108 95 50 14 4 0 0 0 evil_myc11x11 121 78 44 23 10 6 4 0 0 evil_s3m25x5 125 59 51 45 22 10 4 0 2 hamming6_2 64 0 0 0 0 0 0 0 0 hamming6_4 64 15 14 16 10 10 11 10 9 johnson8_2_4 28 14 12 10 9 0 0 0 0 johnson8_4_4 70 29 24 0 0 0 0 0 0 johnson16_2_4 120 31 25 8 19 12 8 4 0 MANNa_9 45 45 43 13 6 0 0 0 0 paley61 61 20 15 15 12 12 10 8 9 paley73 73 19 20 16 14 11 11 10 9 paley89 89 20 22 22 19 16 16 9 14 paley97 97 18 22 17 12 18 12 11 16 paley101 101 26 18 22 19 12 18 14 0
Data presented in Table 3 indicate that when the calculations are performed with $\beta\in\{\frac{1}{10},\frac{1}{8}\}$, the components we need to consider are quite large compared to the size of the graph. For example, for the instance MANN_a9, the components encompassed nearly all vertices of the graph. The largest components contained $50$ vertices when the calculations were executed with $\beta=\frac{1}{6}$. For all other values of the parameters, we observe that the component sizes were relatively small, suggesting that the recomputation can be carried out efficiently.
@@ 6.2.2 Results for larger instances
Instance $n$ $\beta=1/10$ $\beta=1/8$ $\beta=1/6$ $\beta=1/4$ $\beta=1/2$ $\beta=1$ $\beta=10$ $\beta=100$ brock800_1 800 38 33 18 3 0 0 0 0 brock800_2 800 38 35 16 2 0 0 0 0 brock800_3 800 38 31 8 5 0 0 0 0 brock800_4 800 37 32 17 0 0 0 0 0 p_hat1500_1 500 22 20 14 3 0 0 0 0 p_hat1500_2 1500 112 95 34 11 0 0 0 0 p_hat1500_3 1500 156 128 34 7 0 0 0 0
@@ 7 Conclusion
Since not all instances can be embedded on the D-Wave’s QPU, we explored methods to address this issue. We introduced the simple CH-partitioning method as an alternative to the D-Wave’s hybrid solver for certain instances. In this approach, we break down the original problem into smaller subproblems, solve them using the QPU, and then take the largest stable set across the subproblems as the final solution. To improve the quality of results, we combined the proposed partitioning method with the post-processing procedure and obtained promising results. Future research should concentrate on creating techniques to break down partitions further, recognizing partitions that could lead to optimal solutions, and investigating additional efficient reduction strategies to decrease computational costs.
Finally, in order to assess the impact of different $\beta$ values on the quality of solutions, we conducted thorough numerical experiments with various values of the penalty term $\beta$ ranging from $\frac{1}{10}$ to $100$. Our computational results indicated that $\beta=\frac{1}{2}$ produced the best quality solutions, which is in accordance with the given theoretical guarantees.
@@
The program code associated with this paper is available as ancillary files from the arXiv page of this paper. Additionally, the source code and data are available upon request from the authors.

### v2

@@ 4.2 Analysis of samples
Let us return to Example 6. For a subset of vertices $X\subseteq V$ with $|X|=30$ and $|E(G[X])|=5$, Proposition 5 assures us that we can always extract a stable set of size at least $25$ from it. However, it may occur, for instance, that one of the vertices has a degree of $5$, and by removing it, we obtain a stable set of size $29$. This property is generalized with the annihilation number of the graph, a graph parameter that is also an uspper bound on the stability number of a graph, as shown in [30].
@@
The original paper details the justification for this upper bound, but the concept is rather straightforward. Let degrees of vertices be defined as above. In each iteration, we remove one vertex, and we repeat this until we have a stable set. To accelerate the formation of a stable set, we aim to remove the maximum number of edges in each iteration. Consequently, we prioritize removing the vertex with the highest degree in the graph, specifically vertex $v_{n}$. Removing the vertex of degree $d_{n}$, the sum of all degrees drops by $2d_{n}$. Here we assume the best case scenario, which is that this reduction occurs in vertices $v_{i},\ldots,v_{i+d_{n}-1}$, where $v_{i}$ represents the first non-zero degree vertex. By preserving the highest degree vertices for as long as possible, we accelerate the attainment of a stable set. When we remove enough vertices, the total number of edges will drop to zero, leaving the remaining vertices as the upper bound for the stability number of $G$.
@@
Altogether, combining Proposition 5, Theorem 9, and Observation 10, we develop the method for enhancing the solutions returned by a quantum annealer. In the following text, we provide a detailed explanation of this method, which we call the post-processing procedure.
@@ 4.3 Post-processing procedure
Let $S=S(G,\beta)$ denote the solver we use to obtain the initial array of solutions $[X]=[X_{1},\ldots,X_{k}]$, and let $S_{post}=S(G[X_{i}],\beta_{post})$ denote the solver for recomputing the stability number of a sample $X_{i}$. The post-processing steps are as follows: first, we take a look at $X_{1}$, which has the lowest cost, and define the current best solution as $\operatorname{\textit{best}}=|X_{1}|-|E(G[X_{1}])|$. If $X_{1}$ is a stable set, then $|E(G[X_{1}])|=0$, and $|X_{1}|=-\text{cost}(X_{1},\beta)$. Otherwise, from Proposition 5, we know that we can extract a stable set of size at least $|X_{1}|-|E(G[X_{1}])|$ from $X_{1}$.
@@ 5 Dealing with large instances
As mentioned in Section 1.2, some graphs are too large to be directly embedded on the QPU. In such cases, we can employ the method of CH-partitioning and divide a graph into multiple smaller subgraphs, which can then be embedded on the QPU. This method considers a related problem, namely the maximum clique problem, but, as we will observe, the conversion between the maximum clique and the maximum stable set is trivial. Therefore, we can use the given method within our framework for finding the maximum stable set. In Section 5.1, we describe the CH-partitioning method, and in Section 5.2, we outline a modified version of the CH-partitioning we will use in our work. Furthermore, in Section 5.3, we discuss bounding procedures that enable us to reduce the number of partitions we have to consider and present an algorithmic framework that combines the proposed modified partitioning with the post-processing procedure.
@@ 5.1 CH-partitioning
Assume we use the presented CH-partitioning method in the following way. Let $G=(V,E)$ be a graph, where the vertices $V$ are ordered in some ordering, and set $C_{i}=\{v_{i}\}$ for all $i\in\{1,\ldots,n\}$. Then $H_{i}=N(v_{i})\setminus\{v_{i}\}$ for all $i\in\{1,\ldots,n\}$. Altogether, this CH-partition consists of $n$ pairs $(C_{i},H_{i})$ and its cost is $\Delta(G)+1$. This is the CH-partition with the lowest cost function if we enumerate over all $s$ since each partition has to contain in each set $C_{i}$ at least one vertex and all of the neighbors of $C_{i}$.
An advantage of this approach is that the computation of the CH-partitioning itself is not computationally intensive. However, this method also has a disadvantage. Let us assume that $K$ is a maximum clique such that $|K|=k$. Then, $k$ partitions contain the same maximum clique. Thus, we perform computations on $k$ sets containing the same maximum clique. To overcome this obstacle, we use a modified version of the CH-partitioning.
@@ 5.2 Simple CH-partitioning
Our goal is to develop a procedure that reduces the number of partitions containing a maximum clique. To achieve this, we build upon the concept introduced in [39]. In that work, the authors explored upper bounds on clique numbers by disregarding subgraphs. We have now adapted this concept to our CH-partitioning framework. With this adaptation, we define the simple CH-partitioning as follows.
@@
Finally, we note that using simple partitioning greatly reduces the cost of partitioning. We consider some instances from the literature that cannot be directly embedded on QPU and compare the costs of regular CH-partitioning for $s=n$ with the costs for the proposed modified version, where we sort the vertices based on their degrees in increasing order. The results are presented in Table 1.
@@ 5.3 Bounding the partitions
Furthermore, we again use the annihilation number, as defined in Definition 8. Recall that the annihilation number $a(G)$ of a graph $G$ is the largest integer $k$ such that there exist $k$ different vertices in $G$ with the degree sum at most $m$, and represents an upper bound on the stability number of $G$, see Theorem 9. Thus, the annihilation number can be used as a reduction strategy: if the annihilation number of a partition is less than or equal to the size of the largest stable set found in other partitions, that partition can be disregarded. We also apply post-processing to every solution, removing excess edges and potentially further improving the solution.
@@
In our computational implementation of Algorithm 2, we use solvers $S$, $S_{post}$, and penalty terms $\beta$, $\beta_{post}$ as described in Section 4.3, and we sort the vertices based on their degrees in increasing order.
However, it is important to note that alternative vertex orderings could also be employed. Specifically, while introducing the concept of estimating clique size by discarding subgraphs in [39], the authors explored different vertex orderings and found that these can lead to significantly different upper bounds. Therefore, investigating the impact of vertex ordering within the context of simple CH-partitioning combined with post-processing remains a topic for future research.
@@ 6 Computational results
All computations presented in this work were conducted on an Intel(R) Core(TM) i7-9750H with 16GB of RAM. Additionally, we used two cloud solvers from the D-Wave Leap platform: the QPU solver Advantage_system4.1 and the hybrid solver hybrid_binary_quadratic_model_version2. We also utilized a classical, locally run simulated annealing solver, which is part of the D-Wave’s Python library we employed for interacting with the QPU. Finally, we note that the computations presented in this paper were conducted at the end of 2023 and the beginning of 2024. For these computations, the only hyperparameter we adjusted in the D-Wave solvers was the number of runs; all other hyperparameters were left at their default settings.
@@ 6.2.1 Results for smaller instances
A potential drawback of our post-processing method when using $\beta<\frac{1}{2}$ is the increased computational expense compared to $\beta=\frac{1}{2}$ and $\beta>\frac{1}{2}$. This higher cost arises from the need to process a larger quantity of samples. For example, with C125.9 and $\beta=\frac{1}{10}$, we had to recompute stability numbers on all $1000$ samples, while with $\beta=\frac{1}{2}$, recomputation was needed on only one sample. Additionally, the computational effort varies between $\beta=\frac{1}{2}$ and $\beta>\frac{1}{2}$, with $\beta>\frac{1}{2}$ generally requiring less effort, as seen in instances like dsjc125.5. However, this trend reverses for instances such as paley89 and paley101, where $\beta=1$ required more recalculations than $\beta=\frac{1}{2}$.
Recall that we recalculate the stability numbers of samples by considering their connected components. To explore the structure of the samples and evaluate the complexity of these recalculations, the cardinalities of the largest components considered are presented in Table 3.
@@
Instance $n$ $\beta=1/10$ $\beta=1/8$ $\beta=1/6$ $\beta=1/4$ $\beta=1/2$ $\beta=1$ $\beta=10$ $\beta=100$ C125.9 125 67 60 47 18 3 1 1 2 dsjc125.5 125 26 26 19 19 21 14 1 15 dsjc125.9 125 66 59 45 24 1 1 1 1 evil_chv12x10 120 71 50 31 12 6 4 1 1 evil_myc5x24 120 108 95 50 14 4 1 1 1 evil_myc11x11 121 78 44 23 10 6 4 1 1 evil_s3m25x5 125 59 51 45 22 10 4 1 2 hamming6_2 64 1 1 1 1 1 1 1 1 hamming6_4 64 15 14 16 10 10 11 10 9 johnson8_2_4 28 14 12 10 9 1 1 1 1 johnson8_4_4 70 29 24 1 1 1 1 1 1 johnson16_2_4 120 31 25 8 19 12 8 4 1 MANNa_9 45 45 43 13 6 1 1 1 1 paley61 61 20 15 15 12 12 10 8 9 paley73 73 19 20 16 14 11 11 10 9 paley89 89 20 22 22 19 16 16 9 14 paley97 97 18 22 17 12 18 12 11 16 paley101 101 26 18 22 19 12 18 14 1
Data presented in Table 3 indicate that when the calculations are performed with $\beta\in\{\frac{1}{10},\frac{1}{8}\}$, the components we need to consider are quite large compared to the size of the graph. For example, for the instance MANN_a9, the components encompassed nearly all vertices of the graph. The largest components contained $13$ vertices when the calculations were executed with $\beta=\frac{1}{6}$. For all other values of the parameters, we observe that the component sizes were relatively small, suggesting that the recomputation can be carried out efficiently.
@@ 6.2.2 Results for larger instances
Instance $n$ $\beta=1/10$ $\beta=1/8$ $\beta=1/6$ $\beta=1/4$ $\beta=1/2$ $\beta=1$ $\beta=10$ $\beta=100$ brock800_1 800 38 33 18 3 1 1 1 1 brock800_2 800 38 35 16 2 1 1 1 1 brock800_3 800 38 31 8 5 1 1 1 1 brock800_4 800 37 32 17 1 1 1 1 1 p_hat1500_1 500 22 20 14 3 1 1 1 1 p_hat1500_2 1500 112 95 34 11 1 1 1 1 p_hat1500_3 1500 156 128 34 7 1 1 1 1
@@ 7 Conclusion
Since not all instances can be embedded on the D-Wave’s QPU, we explored methods to address this issue. We used the simple CH-partitioning method as an alternative to the D-Wave’s hybrid solver for certain instances. In this approach, we break down the original problem into smaller subproblems, solve them using the QPU, and then take the largest stable set across the subproblems as the final solution. To improve the quality of results, we combined the proposed partitioning method with the post-processing procedure and obtained promising results. Future research should concentrate on creating techniques to break down partitions further, recognizing partitions that could lead to optimal solutions, and investigating additional efficient reduction strategies to decrease computational costs. In particular, it would be valuable to investigate the impact of different vertex orderings.
Furthermore, in order to assess the impact of different $\beta$ values on the quality of solutions, we conducted thorough numerical experiments with various values of the penalty term $\beta$ ranging from $\frac{1}{10}$ to $100$. Our computational results indicated that $\beta=\frac{1}{2}$ produced the best quality solutions, which is in accordance with the given theoretical guarantees. Nevertheless, it would be valuable to conduct further computational experiments by exploring values of $\beta$ in the form $\beta=\frac{1}{2}\pm\epsilon$, where $\epsilon$ is small.
Finally, the only hyperparameter we adjusted in the D-Wave solvers for the computations presented in this work was the number of runs, while all other hyperparameters were left at their default settings. Future research should explore varying additional parameters of the QPU, such as annealing time, spin reversal transforms, and chain strength. Optimizing these parameters could enhance the performance of the QPU before applying the post-processing procedure.
@@
The program code associated with this paper is available as ancillary files from the arXiv page of this paper: arXiv:2405.12845. Additionally, the source code and data are available upon request from the authors.